OpenStudy (anonymous):

Please help waiting for hours for answer....... If you are kind enough to help me!

OpenStudy (anonymous):

The hypotenuse of a right triangle is √18 units. What are the lengths of the legs of the triangle? Will fan and medal!!!!!!!!!!!!!!!

OpenStudy (anonymous):

Can u atleast help me find 3 different answers and explain your strategies/ show work..........................

OpenStudy (welshfella):

This is done by the Pythagoras theorem c^2 = a^2 + b^2 where c = hypotenuse and a and b are the legs

OpenStudy (welshfella):

there are many solutions to your question if the legs of the triangle are equal then c^2 = 2 a^2 (√18 )^2 = 2 a^2 18 = 2 a^2 a^2 = 9 a = 3 - the length of each leg is 3

OpenStudy (anonymous):

Its a right angle triangle

OpenStudy (welshfella):

Yes - thats why you use the Pythagoras theorem - which is true for all right angled triangles

OpenStudy (welshfella):

- there is one triangle which is right angled and isocseles

OpenStudy (anonymous):

The length for each leg is going to be different

OpenStudy (welshfella):

- no not this one:-